iuytyuioiuytyuiop:

There are two numbers which sum to 23. If one number is seven less than four times the other, find the numbers.

iuytyuioiuytyuiop:

@AZ

AZ:

So let's call one number 'x' and the other number 'y' Does that sound good to you?

iuytyuioiuytyuiop:

yep

AZ:

so it says the two numbers have a sum of 23 that means when we add the two numbers we get 23 how can we write that with x and y

iuytyuioiuytyuiop:

x + y = 23

AZ:

Excellent!

AZ:

now we have to write an equation for the other sentence as well one number is seven less than four times the other the way I like to think about it is change the word `is` to an `=` sign when we say less than, what we doing? is that addition/subtraction/multiplication/division

iuytyuioiuytyuiop:

subtraction

iuytyuioiuytyuiop:

x = 7 - 4 * y

AZ:

you almost had it but when we say 7 less than (something) that means it's (something) - 7

AZ:

so x = 4y - 7 does that make sense?

iuytyuioiuytyuiop:

yep

AZ:

so we have two equations now x + y = 23 x = 4y - 7 let me add colors so you can see how you can plug one equation into the other

AZ:

\(\color{red}{x} + y = 23\) \(\color{red}{x} = \color{red}{4y - 7} \) so do you see how we have an x in the first equation? and in the second equation, we say x = something if x is equal to that, then that means we can replace x with that something so you can plug that red part into the first equation

AZ:

 do you see how can you replace it? since they're equal

iuytyuioiuytyuiop:

yep

iuytyuioiuytyuiop:

whats the next step

AZ:

just solve it 4y -7 + y = 23 solve for y

iuytyuioiuytyuiop:

-3y+y=23

AZ:

no no no how'd you get that you have 4 apples and you add another apple, how many apples? think of the y as an apple 4y + y = ??

iuytyuioiuytyuiop:

6y

AZ:

are you sure? you have 4 apples and I give you 1 more apple. How many apples do you have now?

iuytyuioiuytyuiop:

5y

AZ:

there we go!

AZ:

so 4y -7 + y = 23 now we get 5y - 7 = 23 add 7 on both sides, what do you get

iuytyuioiuytyuiop:

12y=30

AZ:

Remember 5y is our apples 7 is just a number Can you add apples and numbers together? No we can't. So when we add 7 on both sides 5y - 7 + 7 = 23 + 7 -7 + 7 gets us 0 and so we're left with 5y = 30

AZ:

so can you solve for y now?

iuytyuioiuytyuiop:

5 * 6 = 30

AZ:

so what is y

iuytyuioiuytyuiop:

y=6

AZ:

Good! Now let's find x and we'll be all done remember we said that the two numbers add up to 23 if one number is 6, what's the other number? what number plus 6 gives us 23

iuytyuioiuytyuiop:

17+6=23

AZ:

so what are you two numbers going to be?

iuytyuioiuytyuiop:

6,17

AZ:

and that's your answer!

iuytyuioiuytyuiop:

thanks

AZ:

You're welcome!